I need the derivative of : y = sqrt(x^2) + 1 * sqrt(x^2) - 1
The answer is supposed to be y' = 2(x^3)((x^4-1)^-1/2)
Help me please.

Answer by Alan3354(69443)

I need the derivative of : y = sqrt(x^2) + 1 * sqrt(x^2) - 1
------------
I saw this one yesterday.
It still makes no sense.
-----
y = sqrt(x^2) + 1 * sqrt(x^2) - 1
y = x + 1*x - 1
y = 2x - 1
y' = 2